OpenStudy (anonymous):

use quadratic formula 2x^2-3x=-10

OpenStudy (anonymous):

2x^2 - 3x + 10 = 0 a = 2 b = -3 c = 10 Use the formula.

OpenStudy (anonymous):

i have so far and i dont know if this is correct 8-27+10

OpenStudy (anonymous):

This problem does not have real roots.

OpenStudy (anonymous):

Determinant = 3^2 - 4*2*10 = 9-80 = -71.

OpenStudy (anonymous):

i have to use a solution set . i dont have any complex numbers in terms of (i)?

OpenStudy (amistre64):

no real roots, means that the roots are complex

OpenStudy (amistre64):

since completing the square is proof of the quad formula: 2x^2 -3x = -10 x^2 -(3/2)x = -5 x^2 -(3/2)x+(3/4)^2 = -5+(3/4)^2 (x-(3/4))^2 = -5+(3/4)^2 x-(3/4) = +- sqrt(-5+(3/4)^2) x = (3/4) +- sqrt(-5+(3/4)^2)

OpenStudy (amistre64):

the wolf compacts that to: \[\frac{3\pm i \sqrt(71)}{4}\]

OpenStudy (anonymous):

is that the answer? ive use the (i) in answers but this one has totally lost me